Find ∫ ln 3π‘₯ βˆ’ 1 π‘₯ 5 + cos 4π‘₯ βˆ’ 𝑒 βˆ’π‘₯ 2 ⁄ 𝑑x


1
Expert's answer
2021-06-07T18:39:03-0400

"\int ln(3x) -x^5 +cos(4x) -e^{\dfrac{-x}{2}}"


"\int 1\times ln(3x) -x^5 +cos(4x) -e^{\dfrac{-x}{2}}"


"[ln(3x)x-\int x \times \dfrac{1}{x} ] +\dfrac{sin(4x)}{4}-{ \dfrac{x^6}{6} } -(-2e^{\dfrac{-x}{2}})"


"xln(3x)- x +\dfrac{sin(4x)}{4}-{ \dfrac{x^6}{6} } +2e^{\dfrac{-x}{2}} +C"